# DeTPP: Leveraging Object Detection for Robust Long-Horizon Event Prediction

### v1

@@ 5 Experiments
We conducted a series of experiments using the HoTPP benchmark (Karpukhin et al., 2024) to evaluate the performance of DeTPP against several popular MTPP approaches. Specifically, we compare DeTPP to RMTPP (Du et al., 2016), NHP (Mei & Eisner, 2017), ODE-RNN (Rubanova et al., 2019). Additionally, we implement a simple baseline, MAE-CE, which is trained with a combination of Mean Absolute Error (MAE) and cross-entropy losses. This baseline is similar to a simplified version of the Intensity-free approach (Shchur et al., 2019). To advance our analysis, we extend the MAE-CE method to predict multiple future events in parallel, referring to the resulting method as MAE-CE-K. The datasets used in this study include Retweet (Zhao et al., 2015), Amazon (Jianmo, 2018), and StackOverflow (Jure, 2014).
5.1 Events Forecasting
We evaluate both next-event and long-horizon metrics. For the next event we measure label prediction accuracy and time step Mean Average Error (MAE). We evaluate long-horizon prediction by Optimal Transport Distance (OTD) and Temporal mAP (T-mAP). The results from Table 1 show that DeTPP may not predict the next event as accurately as the best-performing models, but it significantly outperforms them in long-horizon predictions. The high T-mAP scores achieved by DeTPP can be attributed to our training objective, which resembles the T-mAP metric. Interestingly, DeTPP also improves the optimal transport distance (OTD) metric, suggesting that our training procedure enhances overall model stability rather than merely optimizing for a specific metric.
@@ 5.2 Predictions diversity
As we qualitatively demonstrated in Section 3, autoregressive methods tend to produce repetitive outputs. Here, we provide additional quantitative results that further highlight the low output diversity of traditional approaches. In Table 1, we report the entropy of predicted labels and the differential entropy of predicted time deltas. The results show that in most cases baseline autoregressive methods exhibit low entropy in both predicted labels and time deltas, indicating limited diversity. In contrast, DeTPP usually generate sequences with higher entropy values, reflecting greater diversity in their predictions. It’s important to note that we did not tune the sampling temperature hyperparameter, as increasing temperature typically leads to a reduction in prediction quality.
Table 2: Training epoch and validation times (in seconds) for each method.
Method Retweet Amazon StackOverflow
MAE-CE 5.2 1.8 2.1 1.2 1.1 0.8
RMTPP 4.7 10.2 2.2 5.3 1.1 2.9
NHP 36.5 66.7 12.4 40.0 4.1 19.0
ODE 144.3 145.5 42.2 100.0 12.7 50.0
MAE-CE-K 5.4 0.9 2.3 0.8 1.1 0.6
DeTPP 10.8 1.7 3.5 1.1 1.6 0.9
@@ 5.3 Training and Inference Speed
An important practical aspect of the model is its computational efficiency. Table 2 presents a comparison of the training and inference times for each method. As shown, the training time for DeTPP is approximately twice as long as that of the discrete-time models (MAE-CE, RMTPP, MAE-CE-K). However, DeTPP significantly outperforms continuous-time models (NHP, ODE-RNN). In terms of inference speed, DeTPP is comparable to MAE-CE, slightly slower than MAE-CE-K, but considerably faster than the other methods. This makes DeTPP one of the fastest methods for inference.
5.4 Ablations
We conducted ablation studies to compare DeTPP with the Next-K variant of the MAE-CE method. The results are summarized in Table 1. MAE-CE-K predicts $\mathrm{K}$ future events and is trained by computing pairwise losses between predictions and targets at corresponding positions, bypassing the matching stage. Our findings indicate that while DeTPP exhibits a higher next-item error, it significantly outperforms the Next-K approach in long-horizon prediction tasks. We conclude that the primary advantage of DeTPP lies in its matching-based training loss and specialized inference procedure, rather than just the parallel prediction of multiple future events.
@@ 6 Conclusion
In this work, we introduced DeTPP, a novel event prediction method that leverages concepts from object detection to address the challenges of long-horizon prediction. Our experiments demonstrate that DeTPP effectively overcomes the limitations of traditional autoregressive approaches. Notably, DeTPP generates more diverse predictions, significantly enhancing long-horizon prediction accuracy. Additionally, DeTPP is computationally more efficient, as it predicts multiple future events in parallel. This approach paves the way for advancing the modeling of Marked Temporal Point Processes, offering new opportunities across a wide range of practical applications.

### v2

@@ 5 Experiments
We conducted a series of experiments using the HoTPP benchmark (Karpukhin et al., 2024) to assess the performance of DeTPP against several widely-used MTPP models. Specifically, we compare DeTPP to Intensity-Free (Shchur et al., 2019), intensity-based RMTPP (Du et al., 2016) and NHP (Mei & Eisner, 2017) approaches, ODE-RNN (Rubanova et al., 2019), the transformer-based AttNHP model Yang et al. (2022), and the long-horizon rescoring method HYPRO (Xue et al., 2022). The datasets employed in this study include Retweet (Zhao et al., 2015), Amazon (Jianmo, 2018), StackOverflow (Jure, 2014), MIMIC-IV (Johnson et al., 2023), and Transactions (AI-Academy for teens, 2021), which represent a diverse range of domains and scales, as summarized in Table 1.
Table 1: Datasets statistics and evaluation parameters
Dataset Domain Sequences Events Classes Time unit OTD steps Horizon / Mean length StackOverflow Social. net. 2k 138k 22 Minute 10 10 / 12.0 Amazon Social. net. 9k 403k 16 N/A 5 10 / 14.8 Retweet Social. net. 23k 1.3M 3 Second 10 180 / 14.7 MIMIC-IV Medical 120k 2.4M 34 Day 5 28 / 6.6 Transactions Financial 50k 43.7M 203 Day 5 7 / 9.0
5.1 Long-Horizon Events Forecasting
We evaluate long-horizon prediction performance using Optimal Transport Distance (OTD) and Temporal mean Average Precision (T-mAP). As shown in Table 2, DeTPP significantly outperforms popular MTPP approaches, with the only exception being the OTD metric on the MIMIC-IV dataset, where IFTPP ranks first and DeTPP second. In Appendix C, we show that IFTPP’s lower error stems from a large number of zero time steps in MIMIC-IV. In the latter case, precise events ordering becomes more important than the accurate time modeling. Overall, DeTPP achieves state-of-the-art performance in 9 out of 10 comparisons. The high T-mAP scores of DeTPP can be linked to its training objective, which utilizes matching, similar to T-mAP. However, DeTPP consistently improves the OTD metric, suggesting that its training process enhances overall model performance rather than merely optimizing for a single evaluation criterion.
Table 2: Evaluation results in the long-horizon prediction task. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
Model Metrics (OTD / T-mAP) StackOverflow Amazon Retweet MIMIC-IV Transactions IFTPP 13.64 / 8.31% 6.52 / 22.56% 172.7 / 31.75% 11.53 / 21.67% 6.90 / 5.88% $\pm$0.05 / $\pm$0.50$\pm$0.05 / $\pm$0.52$\pm$4.4 / $\pm$4.44$\pm$0.01 / $\pm$0.21$\pm$0.01 / $\pm$0.13RMTPP 13.17 / 12.72% 6.57 / 20.06% 166.7 / 44.74% 13.71 / 21.08% 6.88 / 6.69% $\pm$0.05 / $\pm$0.16$\pm$0.03 / $\pm$0.33$\pm$3.3 / $\pm$0.94$\pm$0.03 / $\pm$0.29$\pm$0.01 / $\pm$0.12NHP 13.24 / 11.96% 9.02 / 26.29% 165.8 / 45.07% 18.60 / 7.32% 6.98 / 5.61% $\pm$0.02 / $\pm$0.40$\pm$0.35 / $\pm$0.55$\pm$1.6 / $\pm$0.34$\pm$0.19 / $\pm$1.33$\pm$0.01 / $\pm$0.05AttNHP 13.30 / 11.13% 7.30 / 14.62% 171.6 / 25.85% 14.68 / 22.46% 7.50 / 1.48% $\pm$0.02 / $\pm$0.32$\pm$0.06 / $\pm$0.80$\pm$1.0 / $\pm$1.08$\pm$0.08 / $\pm$0.40N/A / N/A ODE 13.27 / 10.52% 9.46 / 22.96% 165.3 / 44.81% 14.74 / 15.18% 6.97 / 5.52% $\pm$0.03 / $\pm$0.23$\pm$0.08 / $\pm$0.61$\pm$0.5 / $\pm$0.69$\pm$0.34 / $\pm$0.15$\pm$0.01 / $\pm$0.13HYPRO 13.26 / 14.69% 6.61 / 20.53% 170.7 / 46.99% 14.87 / 16.77% 7.05 / 7.05% N/A / N/A N/A / N/A N/A / N/A N/A / N/A N/A / N/A DeTPP 12.06 / 23.11% 5.98 / 37.18% 134.4 / 57.37% 12.85 / 30.63% 6.66 / 9.17% $\pm$0.04 / $\pm$0.08$\pm$0.01 / $\pm$0.07$\pm$0.8 / $\pm$0.67$\pm$0.26 / $\pm$0.14$\pm$0.03 / $\pm$0.11%
5.2 Next Event Prediction
Marked temporal point processes are usually evaluated based on the quality of next-event prediction. We measured the next-event type error rate and mean absolute time error across various datasets, with the results presented in Fig. 4. It was observed that DeTPP struggles to accurately model the next event. To address this, we introduced a modified version of DeTPP, referred to as DeTPP+. DeTPP+ incorporates the loss function from the IFTPP method into the first output head, alongside the detection loss, forcing the model to handle both long-horizon and next-event prediction tasks simultaneously. This modification not only brings the model’s performance in line with the best-performing methods but also reduces error on the Transactions dataset, the only dataset where the difference between top methods is significant. This improvement highlights DeTPP’s potential to scale effectively for more complex problems involving a large number of classes and events.
Figure 4: The next event prediction error.
@@ 5.3 Predictions diversity
As we qualitatively demonstrated in Section 3, autoregressive methods tend to produce repetitive outputs. Here, we provide additional quantitative results that further highlight the low output diversity of traditional approaches.
A popular way to make diverse predictions is to adjust temperature parameter during sampling. When temperature approaches zero, the model predicts the label with the maximum output probability, as was done in previous experiments. When temperature becomes large, the model samples from the uniform labels distribution, leading to the highest possible entropy. We therefore need to study the impact of the temperature on both the predictions diversity and long-horizon quality. We do this by varying temperature from 0 to 10 and measuring the average entropy of predicted event types within the horizon. We also evaluate the long-horizon prediction quality in terms of OTD, as T-mAP doesn’t depend on temperature. The results from Fig. 5 show, that DeTPP achieves the optimal balance between diversity and prediction quality on 4 out of 5 datasets. Some methods can be improved by using sampling instead of maximum probability prediction, but they still can’t reach the prediction quality of DeTPP.
Figure 5: The relation between OTD and predictions diversity for varying sampling temperature values. The optimal quality corresponds to the bottom right corner, combining low error and high diversity.
@@ 5.4 Training and Inference Speed
A key practical aspect of the model is its computational efficiency. Fig. 6 presents a comparison of the training and inference times in terms of Requests Per Second (RPS), i.e. the number of processed batch elements [^2] The experiment was done with a single Nvidia RTX 4060 GPU.. The results show that DeTPP has a moderate training time while being the fastest method during inference across all datasets except Transactions, where it is slightly slower than IFTPP due to the computational cost associated with the prediction head. We therefore conclude, that DeTPP is among the most computationally efficient methods in the field.
Figure 6: Computation speed measured as Requests Per Second (RPS) during training and sequence generation.
5.5 Heads Specialization
To provide greater intuition on the functioning of DeTPP, we conducted an additional analysis of the outputs generated by the model’s prediction heads. During training, each prediction head is matched to its closest corresponding ground truth event, leading to a gradual and persistent specialization. This specialization is illustrated in Fig. 7, where we present per-head output statistics. As shown, each prediction head becomes responsible for a specific time interval and a subset of event classes. Notably, the time intervals tend to be short in relation to the overall horizon length, while the number of assigned event classes can vary from 1 to 10. From this, we infer that the specialization of the prediction heads is driven more by temporal factors than by specific event types.
In the Retweet dataset, approximately half of the heads focus on events occurring in the near future, leaving the later parts of the horizon largely uncovered. This suggests that the model naturally excludes rare and distant events, which are challenging to predict accurately. This behavior highlights DeTPP’s inherent flexibility in adapting to the underlying data distribution without additional modifications.
Figure 7: Output statistics across $K$ model heads. Head indices are sorted according to the mean predicted timestamp. The number of classes for Transactions is limited to 20 for clarity.
5.6 Ablation Studies
5.6.1 Comparison with the next-K approach
DeTPP involves two modifications compared to previous methods: a parallel prediction of multiple future events and a novel loss function. To study the impact of these modifications, we evaluated the next-K variant of the IFTPP method. IFTPP-K predicts K future events and applies the IFTPP loss between pairs of predicted and ground truth future events at corresponding positions. As shown in Table 3, IFTPP-K demonstrates a low long-horizon prediction quality, even worse than a simple IFTPP. A possible reason for this failure can be incorrect alignment between predictions and ground truth, leading to noisy gradients and reduced prediction quality. We therefore conclude, that our matching loss is an essential part of the training pipeline.
Table 3: Ablation studies results. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
Model Metrics (OTD / T-mAP) StackOverflow Amazon Retweet MIMIC-IV Transactions IFTPP-K 13.51 / 11.42% 6.68 / 22.57% 167.9 / 34.73% 13.18 / 22.30% 7.19 / 4.43% $\pm$0.06 / $\pm$0.78$\pm$0.01 / $\pm$0.07$\pm$2.8 / $\pm$5.11$\pm$0.05 / $\pm$0.03$\pm$0.00 / $\pm$0.16DeTPP 12.06 / 23.11% 5.98 / 37.18% 134.4 / 57.37% 12.85 / 30.63% 6.66 / 9.17% $\pm$0.04 / $\pm$0.08$\pm$0.01 / $\pm$0.07$\pm$0.8 / $\pm$0.67$\pm$0.26 / $\pm$0.14$\pm$0.03 / $\pm$0.11DeTPP+ 12.14 / 22.72% 5.98 / 37.20% 132.9 / 57.93% 12.95 / 30.35% 6.70 / 9.26% $\pm$0.04 / $\pm$0.32$\pm$0.04 / $\pm$0.06$\pm$0.7 / $\pm$0.33$\pm$0.32 / $\pm$0.25$\pm$0.03 / $\pm$0.09DeTPP-FF 12.27 / 19.87% 5.97 / 37.09% 131.2 / 55.37% 13.32 / 31.13% 6.64 / 9.03% $\pm$0.02 / $\pm$0.26$\pm$0.06 / $\pm$0.09$\pm$0.4 / $\pm$0.40$\pm$0.11 / $\pm$0.17$\pm$0.03 / $\pm$0.23DeTPP-NP 13.62 / 18.08% 6.39 / 36.38% 193.2 / 48.96% 12.52 / 33.62% 6.88 / 9.32% $\pm$0.04 / $\pm$0.27$\pm$0.02 / $\pm$0.25$\pm$1.5 / $\pm$0.28$\pm$0.42 / $\pm$0.10$\pm$0.02 / $\pm$0.07%
5.6.2 DeTPP vs DeTPP+
In Section 5.2, we introduced an extension of DeTPP, called DeTPP+, that can efficiently solve the next event prediction problem. The question may rise, how well the extended method at the long horizon prediction. We compare DeTPP and DeTPP+ in Table 3. It can be seen, that the difference between long-horizon prediction quality of DeTPP and DeTPP+ is not significant. We therefore conclude, that there is no need for two separate models for next-event and long-horizon predictions, and DeTPP+ can be used in both cases.
5.6.3 Conditional head
In Section 4.3, we introduced a conditional head architecture. In Table 3, we present evaluation results for a model with a simple feed-forward head. The resulting approach, called DeTPP-FF, shows slightly better OTD scores and slightly worse T-mAP values. At the same time, DeTPP-FF has much more model parameters. For example, on Retweet it contains 140082 trainable parameters, while DeTPP roughly the half of that size – 71543. We therefore preferred conditional head due to its low memory usage and better scaling properties with the growing number of prediction heads.
5.6.4 Presence matching
One of the fundamental differences between DeTPP and the DeTR method from object detection Carion et al. (2020), is the usage of the presence score for matching, as stated in Equation 6 and Equation 4. The usage of presence score allows the model to select a subset of its heads depending on historical data, better adapting to a particular sequence of events during training. We conducted an additional study and measured the performance of the model with $\mathcal{L_{\mathrm{BCE}}}$ excluded during matching. According to the results from Table 3, the resulting DeTPP-NP is inferior to DeTPP with a large margin on small datasets, while giving moderate or no improvement in other cases. We therefore conclude, that matching with presence score improves the stability of the model among different dataset sizes.
5.6.5 The number of heads
An important hyperparameter of DeTPP is the number $K$ of prediction heads. We analyzed the quality of the model depending on $K$ according to the T-mAP metric, as it, unlike OTD, considers the whole horizon. The results are presented in Fig. 8. It can be seen, that the number of heads is an important hyperparameter, that must be tuned individually for each dataset. In most cases, however, the optimal value lies in an interval 32-64. It is also worth noting that the dependency is not monotonic and the large $K$ might be suboptimal. On the MIMIC-IV and Transactions datasets, the value of $K$ has little impact on performance within a specified range, so we reduced the number of heads to speed up training.
Figure 8: The dependency of T-mAP on the DeTPP hyperparameter $K$. The vertical line indicates the selected value.
6 Limitations and Future Work
While DeTPP effectively addresses both next event and long-horizon prediction tasks, it has certain limitations. First, DeTPP relies on a fixed horizon size, $H$. In this study, we selected $H$ based on the hyperparameters of the OTD and T-mAP metrics. A change in the evaluation metric typically requires adjusting DeTPP’s parameters. A potential direction for future research could explore combining horizon-based and autoregressive inference, enabling DeTPP to make predictions across arbitrary horizons without requiring modifications to the model.
Another limitation of DeTPP is that it independently predicts all events within the horizon. Future work could focus on incorporating rescoring techniques, such as those in Xue et al. (2022), or applying beam search to improve DeTPP’s predictive performance. Additionally, better time modeling might be achieved by integrating intensity-based approaches, like NHP or RMTPP, with DeTPP, offering another promising direction for research.
Furthermore, some techniques from DeTPP could be adapted for object detection in computer vision Carion et al. (2020). DeTPP introduces a probabilistic framework that unifies different loss functions, using the same objective during both matching and backpropagation, which enhances optimization robustness. Notably, DeTPP employs a presence score during matching, which, as shown in our ablation studies, significantly improves the model performance on most datasets.
@@ 7 Conclusion
In this work, we introduced DeTPP, a novel event prediction method that leverages concepts from object detection to address the challenges of long-horizon prediction. Our experiments demonstrate that DeTPP effectively overcomes the limitations of traditional autoregressive approaches. Notably, DeTPP generates more diverse predictions, significantly enhancing long-horizon prediction accuracy. Moreover, DeTPP is computationally more efficient than most MTPP methods, as it predicts multiple future events in parallel. This approach paves the way for advancing the modeling of Marked Temporal Point Processes, offering new opportunities across a wide range of practical applications.
Appendix A On-line Calibration Algorithm
Algorithm 1 On-line calibration
1:batches $b_{n},n=\overline{1,N}$ with size $B$, momentum $m$.
2:prediction thresholds $th_{k},k=\overline{1,K}$.
3:$r_{k}\leftarrow 1,k=\overline{1,K}$                                               $\triangleright$ Initialize head matching rates
4:$th_{k}\leftarrow 0,k=\overline{1,K}$                                             $\triangleright$ and thresholds
5:for $n=1$ to $N$ :                                                $\triangleright$ Process each batch
6:   $\mathrm{matched}_{k}\leftarrow 0,k=\overline{1,K}$
7:   $\mathrm{scores}_{k}\leftarrow\emptyset,k=\overline{1,K}$
8:   for $(y,\hat{y})$ in $b_{n}$ :                                              $\triangleright$ Compute statistics for batch elements
9:      $\sigma\leftarrow$ optimal matching for $(y,\hat{y})$ given by Eq.4.
10:      $\mathrm{matched}_{k}\leftarrow\mathrm{matched}_{k}+\mathrm{~{}I}\left[k\in\sigma\right],k=\overline{1,K}$      $\triangleright$ $\mathrm{I}$ is the indicator function
11:      $\mathrm{scores}_{k}\leftarrow\mathrm{scores}_{k}\bigcup\mathrm{~{}}\{\hat{o}_{k}\},k=\overline{1,K}$
12:   $r_{k}\leftarrow(1-m)r_{k}+m\mathrm{~{}matched}_{k}/B,k=\overline{1,K}$        $\triangleright$ Update matching rates
13:   $th_{k}\leftarrow(1-m)th_{k}+m\mathrm{~{}Q}(r_{k},\mathrm{scores}_{k}),k=\overline{1,K}$      $\triangleright$ and thresholds. $\mathrm{Q}$ is the quantile function
Appendix B Hyperparameters
The number of prediction heads $K$ and matching weights were chosen using a Bayesian optimizer, with the validation T-mAP score as the objective function. The selected horizon aligns with that used in T-mAP. The DeTPP+ loss weight was set to 4 in all cases. The final values are presented in Table 4. Note, that only relative values of matching weights are important. Thus, DeTPP has 3 hyperparameters that must be tuned.
Table 4: DeTPP hyperparameters.
Dataset K Horizon DeTPP+ loss weight Matching weights
Presence Timestamps Labels
Transactions 32 7 4 4 0.47 0.57
MIMIC-IV 32 28 4 4 0.4 0.8
Retweet 64 180 4 4 0.15 0.6
Amazon 48 10 4 4 0.5 0.8
StackOverflow 48 10 4 4 2 1
Appendix C OTD on MIMIC-IV
As discussed in Section 5.1, DeTPP underperforms compared to IFTPP on the MIMIC-IV dataset based on the OTD score. In this section, we analyze a toy example to demonstrate that this behavior is expected and relates to the challenges with time-based ordering present in MIMIC-IV.
First, let’s examine the distribution of time steps in MIMIC-IV. As shown in Fig. 9(a), MIMIC-IV contains a large number of events with near-zero timestamps. At the same time, these events follow a natural order (e.g., admission, laboratory tests, ICU stay start, treatment, ICU stay end). Therefore, in many cases within MIMIC-IV, precise ordering of events is more critical than accurate time prediction.
To illustrate this, consider a simple dataset consisting of a single sequence of 16 events with zero timestamps and labels that form a sequence of integers from 0 to 15. In this dataset, the timestamps provide no information about the event order. However, the OTD score is calculated for sequences of length 5, meaning the model must correctly predict the first 5 events to minimize the score. As shown in Fig. 9(b), only IFTPP and RMTPP achieve a low OTD score, while other methods, including those with high T-mAP scores, perform significantly worse. This outcome is expected for DeTPP, as its ordering relies entirely on the predicted timestamps. Interestingly, the NHP and ODE-based approaches demonstrate similar behavior. This may be due to the NHP objective, which models the arrival times of each event class independently, leading to random ordering when timestamps are identical.
Based on this experiment, we conclude that most TPP methods are not well-suited for order-sensitive tasks on datasets with a high proportion of equal or near-equal timestamps.
(a) MIMIC-IV timestamps distribution.
(b) Prediction quality on the Toy dataset with constant timestamps.
Figure 9: Additional experiments demonstrating specifics of the OTD metric on MIMIC-IV.